General Meeting Summary 11/2/17
CRYSTAL UPDATES (Vladimir, Rong Wang, Tanja)
- Checked subset of boxes of the 320 crystal set for feedback on invoicing - so far all looked ok
- Optical measurements of 320 crystal set ongoing at CUA - currently have 60 crystals (6 boxes)
- Based on microscopy results carried out in collaboration with VSL, crystal quality not as good as Crytur's
- Transmittance tests suggest that most crystals pass requirement at 420nm and 620nm. A large fraction fails the 360nm requirement.
- Light yield tests show a large spread of results from 10 p.e./MeV to almost 30 p.e./MeV, also includes crystals with even higher apparent light yield.
- Sampling method for crystal chemical analysis established in collaboration with VSL
- Looking for correlations between crystals with extreme high/low optical properties and chemical composition
- Discussion about optical measurements methods and future studies
- for accept/reject tests, continue measurements currently ongoing
- to see what can be achieved in terms of resolution will use prototype, perhaps with best/worst set of crystals
- Tests of Crytur crystals from different crystallization number ongoing at CUA
- in collaboration with Giessen Univ.
- also discussions ongoing with Crytur
- Tests of 10 SICCAS crystals ongoing at IPN-Orsay
- observed discrepancies with CUA measurements, but not with as uniform offsets as anticipated
- planning to perform measurements at Giessen Univ. with Valera Dormenev
- Note: recent measurements of Crytur crystals are consistent between CUA and Giessen Univ.
CRYSTAL DATABASE DEVELOPMENT (Vladimir)
- Goal is to have an easy access and search software including all information about crystals and PMTs
- A mysql structure based on that of Hall D is presented - send comments on what to include o Vladimir
- Most recent PMTs have the correct sockets
- Next steps: assembly and mount divider (Fernando)
- PMT testing: Hamlet will check with Fernando about a testing setup at JLab
- Discussion about correlation between measurements with PMTs of different diameters
- Main coil still on schedule for end of year
- Yoke also still on schedule
- The preliminary agenda is presented, fits generally after the Hall C meeting on Tuesday, waiting for Thia for confirmation if NPS meeting can start at 3pm
- TCS talk will be in general Hall C meeting
NEXT MEETING: 16 NOVEMBER at 9AM (EST)